OpenStudy (anonymous):

Find an exact value. cos(-7pi/12)

OpenStudy (anonymous):

Hint: \[Cos(-x)=cosx\] \[Cos(7 pie/12) = Cos (pie - 5pie/12) = -\cos(5pie/12)\] \[Cos(5pie/12) = \cos(75) = \cos(45+30)\] \[Cos(x+y) = cosx.cosy+sinx.siny\]

OpenStudy (anonymous):

thank you